You place 50.0 mL of 0.500 M NaOH in a coffee-cup calorimeter at 35.00ᵒC and add 25.00 mL of
0.500 M HCl, also at 35.00ᵒC. After stirring, the final temperature is 37.21ᵒC. [Assume the total
volume is the sum of the individual volumes and that the final solution has the same density (1.00
g/mL) and specific heat capacity as water (4.184 J/g.K).
(a). Calculate the qsoln (in J and kJ)
(b) Calculate the change in enthalpy, ∆H of the reaction (in kJ/mol of water formed).
1
Expert's answer
2021-10-10T10:18:13-0400

Moles of NaOH"=0.500×\frac{50.0}{1000}=0.025moles"



Total volume "=50.0+25.00=75ml"


Total mass ="75\u00d71.00=75grams"


NaOH (aq) + HCl (aq) → NaCl (aq) + H2O (l)